OpenStudy (anonymous):

Would appreciate some help! :) Give the next two numbers in the pattern. 9, 4, -1, -6,... Write a formula for this sequence: 8, -1, -10, -19,... Write a formula for this sequence: 8, -1, -10, -19,... Write an expression for the sequence: 10, 17, 24, 31,... What is the common difference in this sequence? 10, 17, 24, 31,…

OpenStudy (anonymous):

Give the next two numbers in the pattern. 4, 8, 12, 16,...

OpenStudy (anonymous):

Give the next two numbers in the pattern. 0, 0.75, 1.5, 2.25, ..

OpenStudy (mathmath333):

3.code \(\large\tt \begin{align} \color{black}{9, 4, -1, -6,.....\\~\\ \text{in this sequence the the pattern is showing the}\\ \text{common difference of -5}\\ 4-9=-5\\ -1-4=-5\\ -6-(-1)=-5\\ \text{this is AP sequence} }\end{align}\)

OpenStudy (mrnood):

These are all arithmetic sequences - which means that you add (or subtract) the same number to move to each next step In the first one 9,4,-1,-6 yo u can see that each step adds -5 to the previous. To get the next two steps just carry on from -6, simply add -5, then -5 again... All the others are similar - look for the difference between two adjacent steps - it is the same for all steps in that question

OpenStudy (cwrw238):

for the first sequence :- 5 is subtracted to get next term

OpenStudy (anonymous):

-11, -16 ?

OpenStudy (anonymous):

@cwrw238 @MrNood @mathmath333

OpenStudy (mathmath333):

YES \(\Huge\tt \begin{align} \color{black}{\checkmark }\end{align}\)